# Exponential actions defined by vector configurations, Gale duality, and moment-angle manifolds

### v1

@@ 2. Free orbits (nondegenerate leaves)
$V_{{\textbf{{x}}}}=\{{\textbf{{v}}}\in V\colon\langle\gamma_{i},{\textbf{{v}}}\rangle=0\;\text{ whenever }x_{i}=0\}.$
@@
For the second statement, $V_{{\textbf{{x}}}}\neq 0$ if and only if there is a nonzero ${\textbf{{v}}}\in V_{{\textbf{{x}}}}$ such that $\langle\gamma_{i},{\textbf{{v}}}\rangle=0$ for $x_{i}\neq 0$, which implies that $\{\gamma_{i}\colon x_{i}\neq 0\}\subset\Gamma$ does not span $V^{*}$. ∎
@@ 3. Linear Gale duality
Since $\Gamma$ spans $V^{*}$, the space $W^{*}$ has dimension $n:=m-k$. Consider the dual exact sequence
@@
where the map $\varGamma^{*}$ takes v to $(\langle\gamma_{1},{\textbf{{v}}}\rangle,\ldots,\langle\gamma_{m},{\textbf{{v}}}\rangle)$. Define ${\textbf{{a}}}_{i}=A({\textbf{{e}}}_{i})$. The linear Gale transform takes the vector conguration $\Gamma$ in $V^{*}$ to the vector configuration $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ in $W^{*}$. It is an involutive procedure, so $\mathrm{A}$ is called the Gale dual of $\Gamma$, and $\Gamma$ is the Gale dual of $\mathrm{A}$. Whenever we talk about a pair of Gale dual configurations $\mathrm{A}$ and $\Gamma$, we assume that $\mathrm{A}$ spans $W^{*}$ and $\Gamma$ spans $V^{*}$.
@@
For any $I\subset[m]$, the vectors in $\mathrm{A}_{I}$ are linearly independent in $W^{*}$ if and only if $\Gamma_{\widehat{I}}$ spans $V^{*}$.
@@
4. Polyhedral cones and fans
We collect basic facts about polyhedral cones and fans here, referring to [11], [18, Chapter 1] and [15, Chapters 1 and 3] for a more detailed treatment.
@@ 4. Polyhedral cones and fans
If $\mathcal{C}$ is $\mathrm{A}$-closed, then all faces of a cone in the set $\{\mathop{\mathrm{cone}}\mathrm{A}_{I}\colon I\in\mathcal{C}\}$ also belong to this set. The set $\{\mathop{\mathrm{cone}}\mathrm{A}_{I}\colon I\in\mathcal{C}\}$ is a fan whenever any two cones $\mathop{\mathrm{cone}}\mathrm{A}_{I}$ and $\mathop{\mathrm{cone}}\mathrm{A}_{J}$ intersect in a common face (which has to be $\mathop{\mathrm{cone}}\mathrm{A}_{I\cap J}$). In this case we say that $\{A,\mathcal{C}\}$ is a fan data.
@@
Hence, $\mathop{\mathrm{cone}}\mathrm{A}_{I}$ and $\mathop{\mathrm{cone}}\mathrm{A}_{J}$ are separated by a hyperplane for each $I,J\in\mathcal{C}$, so (a) holds by by Lemma 4.2.
@@
A fan is simplicial if each of its cones is generated by a linearly independent set of vectors. The fan $\Sigma$ defined by a fan data $\{\mathrm{A},\mathcal{C}\}$ is simplicial if and only if $\mathcal{C}$ is a simplicial complex $\mathcal{K}$. That is, a simplicial fan can be written as $\Sigma=\{\mathop{\mathrm{cone}}\mathrm{A}_{I}\colon I\in\mathcal{K}\}$, where $\mathcal{K}$ is a simplicial complex on $[m]$ and $\mathrm{A}_{I}$ is linearly independent for any $I\in\mathcal{K}$. The simplicial complex $\mathcal{K}$ is called the underlying complex of a simplicial fan $\Sigma$.
We say that $\{\mathrm{A},\mathcal{K}\}$ is a simplicial fan data (or a triangulated configuration [7]) if $\Sigma=\{\mathop{\mathrm{cone}}\mathrm{A}_{I}\colon I\in\mathcal{K}\}$ is a fan. This fan is automatically simplicial, since $\mathrm{A}_{I}$ is linear independent for any $I\in\mathcal{K}$. (Indeed, if there is a linear dependency between $a_{i}$, $i\in I$, then there are $J,J^{\prime}\subset I$ such that $\mathop{\mathrm{cone}}\mathrm{A}_{J}$ and $\mathop{\mathrm{cone}}\mathrm{A}_{J^{\prime}}$ overlap.)
A simplicial complex $\mathcal{K}$ defines the coordinate fan $\Sigma_{\mathcal{K}}$ in $\mathbb{R}^{m}$ consisting of coordinate cones $\mathop{\mathrm{cone}}({\textbf{{e}}}_{i}\colon i\in I)$ with $I\in\mathcal{K}$, where ${\textbf{{e}}}_{i}$ is the $i$th coordinate vector. Then $\{\mathcal{K},\mathrm{A}\}$ is a fan data if and only if the restriction of the map $A\colon\mathbb{R}^{m}\to W^{*}$, ${\textbf{{e}}}_{i}\mapsto{\textbf{{a}}}_{i}$, to the union of cones of $\Sigma_{\mathcal{K}}$ is a bijection. In this case, $A(\Sigma_{\mathcal{K}})$ is precisely the fan $\Sigma$ defined by $\{\mathcal{K},\mathrm{A}\}$.
@@
    $\{\mathcal{K},\mathrm{A}\}$ is a simplicial fan data;
@@ 5. Proper actions
We refer to [25, Chapter 21] for the details.
@@
    the action $V\times U(\mathcal{K})\to U(\mathcal{K})$ is proper if and only if $\{\mathcal{K},\mathrm{A}\}$ is a simplicial fan data, in which case the action is also free.
@@
To prove (2), first assume that $\{\mathcal{K},\mathrm{A}\}$ is a fan data. We need to check that the preimage $h^{-1}(C)$ of a compact subset $C\subset V\times U(\mathcal{K})$ under the shear map $h\colon V\times U(\mathcal{K})\to U(\mathcal{K})\times U(\mathcal{K})$ is compact. Since $V$ and $U(\mathcal{K})$ are metric spaces, it suffices to check that any infinite sequence $\{({\textbf{{v}}}^{(k)},{\textbf{{x}}}^{(k)})\colon k=1,2,\ldots\}$ of points in $h^{-1}(C)$ contains a convergent subsequence. Since $C$ is compact, by passing to a subsequence we may assume that the sequence
@@
Now assume that the data $\{\mathcal{K},\mathrm{A}\}$ does not define a fan. Then condition (b) of Theorem 4.4 fails, and we have
@@
Finally, if $\{\mathcal{K},\mathrm{A}\}$ is a simplicial fan data, then $\mathrm{A}_{I}$ is linearly independent for any $I\in\mathcal{K}$, so a proper action $V\times U(\mathcal{K})\to U(\mathcal{K})$ is free by statement (1). ∎
@@ 6. Completeness and compactness
Let $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ be the Gale dual configuration in $V^{*}$. It definin3g action (1.1), whose restriction to $U(\mathcal{K})$ is free and proper by Theorem 5.1. We have the following criterion of compactness for the quotient of the action $V\times U(\mathcal{K})\to U(\mathcal{K})$.
@@
Assume that the data $\{\mathcal{K},\mathrm{A}\}$ defines a simplicial fan $\Sigma$. The quotient smooth manifold $U(\mathcal{K})/V$ is compact if and only if the fan $\Sigma$ is complete.
@@
If $J\neq\varnothing$, then $({\textbf{{v}}}\cdot{\textbf{{x}}}^{(n)})_{j}=0$ for any $j\in J$, ${\textbf{{v}}}\in V$ and $n=1,2,\ldots$ Consider the link subcomplex $\mathop{\mathrm{missing}}{link}_{\mathcal{K}}J=\{I\in\mathcal{K}\colon I\cup J\in\mathcal{K},\,I\cap J=\varnothing\}$. Then the data $\{\mathop{\mathrm{missing}}{link}_{\mathcal{K}}J,{\textbf{{a}}}_{j}\colon j\notin J\}$ defines a complete fan in the quotient space $W^{*}/\langle{\textbf{{a}}}_{j}\colon j\in J\rangle$, which has dimension less than $W^{*}$. Therefore, we can apply the induction assumption in this case.
@@
The condition of completeness (or compactness of the quotient) can be formulated in terms of the fan data $(\mathcal{K},\mathrm{A})$ as follows.
@@
Assume that the data $\{\mathcal{K},\mathrm{A}\}$ defines a simplicial fan $\Sigma$ in an $n$-dimensional space $W^{*}$. Then, the $\Sigma$ is complete if and only if one of the following two equivalent conditions is satisfied:
@@ 7. Real moment-angle manifolds
2. Let $(X,A)=(D^{1},S^{0})$, where $D^{1}$ is a closed interval $[-1,1]$ and $S^{0}$ is its boundary, consisting of two points. The polyhedral product $(D^{1},S^{0})^{\mathcal{K}}$ is known as the real moment-angle complex [12, §3.5][13]:
@@
We have an inclusion of pairs $(D^{1},S^{0})\hookrightarrow(\mathbb{R},\mathbb{R}^{\times})$, which induces an inclusion $\mathcal{R}_{\mathcal{K}}=(D^{1},S^{0})^{\mathcal{K}}\hookrightarrow(\mathbb{R},\mathbb{R}^{\times})^{\mathcal{K}}=U(\mathcal{K})$. So the required homemorphism will follow from the fact that the $V$-orbit of any point ${\textbf{{x}}}=(x_{1},\ldots,x_{m})\in U(\mathcal{K})$ intersects $\mathcal{R}_{\mathcal{K}}$ at a single point.
@@
In [13, Theorem 4.7.5], it is proved that ther is a homotopy equivalence $U(\mathcal{K})\simeq\mathcal{R}_{\mathcal{K}}$ for arbitrary $\mathcal{K}$. By virtue of Theorem 7.3, this homotopy equivalence can be realised as the quotient by a free action of $V\cong\mathbb{R}^{k}$ when $\mathcal{K}$ is the underlying complex of a complete simplicial fan.
@@ 8.1. Normal fans
is a fan in $W^{*}$, referred to as the normal fan of the polyhedron $P$. If $P$ is not full-dimensional, then the cones $\sigma_{Q}$ are not strictly convex, and $\Sigma_{P}$ is a generalised fan [15, Definition 6.2.2]
The support $|\Sigma_{P}|$ of a normal fan is convex (it coincides with the dual of the recession cone of $P$, see [15, Theorem 7.1.6]). In particular, $\Sigma_{P}$ is complete if and only if $P$ is a polytope.
@@
The normal fan $\Sigma_{P}$ of a polyhedron (8.1) consists of cones $\mathop{\mathrm{cone}}\mathrm{A}_{I}$ such that $\sum_{i=1}^{m}b_{i}\gamma_{i}\in\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\Gamma_{\widehat{I}}$.
@@
where the second identity holds because $\Sigma$ has convex support. Since both $\Sigma_{P^{\prime}}$ and $\Sigma$ are subfans of $\Sigma_{P}$, we concude that $\Sigma_{P^{\prime}}=\Sigma$. ∎
@@ 8.2. From polyhedra to quadrics
Now we turn to the quotient (leaf space) $U(\mathcal{K})/V$. As we have seen in Theorem 7.3, the quotient $U(\mathcal{K})/V$ is identified with the real moment-angle manifold $\mathcal{R}_{\mathcal{K}}$ when the data $\{\mathcal{K},{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ defines a complete simplicial fan $\Sigma$. When $\Sigma$ is the normal fan of a polytope, there is another realisation of $U(\mathcal{K})/V$, which has an intrinsic smooth structure as a nondegenerate intersection of real quadratic surfaces.
@@
This result has well-known interpretations in the theory of Hamiltonian torus actions in symplectic geometry and in the theory of toric varieties (see [2, Proposition 3.1.1] and [20, Appendix 1, Theorems 1.2 and 1.4], and also §14.3 below). However, it is purely topological in nature.
@@
It remains to prove the key fact about the image of $L$. The argument below is based on the properties of the Legendre transform and is taken from [20, Appendix 1].
@@
This follows by combining Theorem 7.3 with Theorem 8.4 (4), as both $\mathcal{R}_{\mathcal{K}}$ and the intersection of quadrics are homeomorphic to the quotient $U(\mathcal{K})/V$ (see also [13, Theorem 6.2.4] for a direct argument). Note that the fan is assumed to be complete in Theorem 7.3, so we require that (8.1) is bounded, i. e. a polytope. ∎

### v2

@@ 2. Free orbits (nondegenerate leaves)
$V_{{\textbf{{x}}}}=\{{\textbf{{v}}}\in V\colon\langle\gamma_{i},{\textbf{{v}}}\rangle=0\;\text{ whenever }x_{i}\neq 0\}.$
@@
For the second statement, $V_{{\textbf{{x}}}}\neq 0$ if and only if there is a nonzero ${\textbf{{v}}}\in V_{{\textbf{{x}}}}$ such that $\langle\gamma_{i},{\textbf{{v}}}\rangle=0$ for $x_{i}\neq 0$, which holds precisely when $\{\gamma_{i}\colon x_{i}\neq 0\}\subset\Gamma$ does not span $V^{*}$. ∎
@@ 3. Linear Gale duality
Since $\Gamma$ spans $V^{*}$, the space $W^{*}$ has dimension $n:=m-k$. We identify the dual space of $\mathbb{R}^{m}$ with $\mathbb{R}^{m}$ via the isomorphism that takes the standard basis to its dual basis, and write vectors in $\mathbb{R}^{m}$ by their coordinates in the standard basis. Consider the dual exact sequence
@@
where the map $\varGamma^{*}$ takes v to $(\langle\gamma_{1},{\textbf{{v}}}\rangle,\ldots,\langle\gamma_{m},{\textbf{{v}}}\rangle)$. Define ${\textbf{{a}}}_{i}=A({\textbf{{e}}}_{i})$. The linear Gale transform takes the vector configuration $\Gamma$ in $V^{*}$ to the vector configuration $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ in $W^{*}$. It is an involutive procedure, so $\mathrm{A}$ is called the Gale dual of $\Gamma$, and $\Gamma$ is the Gale dual of $\mathrm{A}$. Then $\mathrm{A}$ spans $W^{*}$ by duality. The linear function ${\textbf{{a}}}_{i}$ takes a vector in $W$ (viewed as a subspace in $\mathbb{R}^{m}$) to its $i$th coefficient.
@@
For any $I\subset[m]$, the vectors in $\mathrm{A}_{I}=\{{\textbf{{a}}}_{i}\colon i\in I\}$ are linearly independent in $W^{*}$ if and only if $\Gamma_{\widehat{I}}=\{\gamma_{j}\colon j\notin I\}$ spans $V^{*}$.
@@
4. Fans and triangulated configurations
@@
If $\mathcal{C}$ is $\mathrm{A}$-closed, then all faces of a cone in the set $\{\mathop{\mathrm{cone}}\mathrm{A}_{I}\colon I\in\mathcal{C}\}$ also belong to this set. The set $\{\mathop{\mathrm{cone}}\mathrm{A}_{I}\colon I\in\mathcal{C}\}$ is a fan whenever any two cones $\mathop{\mathrm{cone}}\mathrm{A}_{I}$ and $\mathop{\mathrm{cone}}\mathrm{A}_{J}$ intersect in a common face (which has to be $\mathop{\mathrm{cone}}\mathrm{A}_{I\cap J}$). In this case we say that $\{\mathcal{C},\mathrm{A}\}$ is a fan data.
@@
Hence, $\mathop{\mathrm{cone}}\mathrm{A}_{I}$ and $\mathop{\mathrm{cone}}\mathrm{A}_{J}$ are separated by a hyperplane for each $I,J\in\mathcal{C}$, so (a) holds by Lemma 4.2.
@@
A fan is simplicial if each of its cones is generated by a linearly independent set of vectors. The fan $\Sigma$ defined by a fan data $\{\mathcal{C},\mathrm{A}\}$ is simplicial if and only if $\mathcal{C}$ is a simplicial complex $\mathcal{K}$. That is, a simplicial fan can be written as $\Sigma=\{\mathop{\mathrm{cone}}\mathrm{A}_{I}\colon I\in\mathcal{K}\}$, where $\mathcal{K}$ is a simplicial complex on $[m]$ and $\mathrm{A}_{I}$ is linearly independent for any $I\in\mathcal{K}$. The simplicial complex $\mathcal{K}$ is called the underlying complex of a simplicial fan $\Sigma$.
A pair $\{\mathcal{K},\mathrm{A}\}$ consisting of a simplicial complex $\mathcal{K}$ on $[m]$ and a spanning vector configuration $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ in $W^{*}$ is called a simplicial fan data or a triangulated configuration [7] if $\Sigma=\{\mathop{\mathrm{cone}}\mathrm{A}_{I}\colon I\in\mathcal{K}\}$ is a fan in $W^{*}$. This fan is automatically simplicial, since $\mathrm{A}_{I}$ is linear independent for any $I\in\mathcal{K}$. (Indeed, if there is a linear dependency between $a_{i}$, $i\in I$, then there are $J,J^{\prime}\subset I$ such that $\mathop{\mathrm{cone}}\mathrm{A}_{J}$ and $\mathop{\mathrm{cone}}\mathrm{A}_{J^{\prime}}$ overlap.)
A simplicial complex $\mathcal{K}$ defines the coordinate fan $\Sigma_{\mathcal{K}}$ in $\mathbb{R}^{m}$ consisting of coordinate cones $\mathop{\mathrm{cone}}({\textbf{{e}}}_{i}\colon i\in I)$ with $I\in\mathcal{K}$, where ${\textbf{{e}}}_{i}$ is the $i$th coordinate vector. Then $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration if and only if the restriction of the map $A\colon\mathbb{R}^{m}\to W^{*}$, ${\textbf{{e}}}_{i}\mapsto{\textbf{{a}}}_{i}$, to the union of cones of $\Sigma_{\mathcal{K}}$ is a bijection. In this case, $A(\Sigma_{\mathcal{K}})$ is precisely the fan $\Sigma$ defined by $\{\mathcal{K},\mathrm{A}\}$.
@@
    $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration;
@@ 5. Proper actions
We refer to [31, Chapter 21] for the details.
@@
    the action $V\times U(\mathcal{K})\to U(\mathcal{K})$ is proper if and only if $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration, in which case the action is also free.
@@
To prove (2), first assume that $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration. We need to check that the preimage $h^{-1}(C)$ of a compact subset $C\subset U(\mathcal{K})\times U(\mathcal{K})$ under the shear map $h\colon V\times U(\mathcal{K})\to U(\mathcal{K})\times U(\mathcal{K})$ is compact. Since $V$ and $U(\mathcal{K})$ are metric spaces, it suffices to check that any infinite sequence $\{({\textbf{{v}}}^{(k)},{\textbf{{x}}}^{(k)})\colon k=1,2,\ldots\}$ of points in $h^{-1}(C)$ contains a convergent subsequence. Since $C$ is compact, by passing to a subsequence we may assume that the sequence
@@
Now assume that the data $\{\mathcal{K},\mathrm{A}\}$ do not define a fan. Then condition (b) of Theorem 4.4 fails, and we have
@@
Finally, if $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration, then $\mathrm{A}_{I}$ is linearly independent for any $I\in\mathcal{K}$, so a proper action $V\times U(\mathcal{K})\to U(\mathcal{K})$ is free by statement (1). ∎
@@ 6. Completeness and compactness
Let $\Gamma=\{\gamma_{1},\ldots,\gamma_{m}\}$ be the Gale dual configuration in $V^{*}$. The restriction of the exponential action (1.1) to $U(\mathcal{K})$ is free and proper by Theorem 5.1. We have the following criterion of compactness for the quotient of the action $V\times U(\mathcal{K})\to U(\mathcal{K})$.
@@
Assume that $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration defining a simplicial fan $\Sigma$. The quotient smooth manifold $U(\mathcal{K})/V$ is compact if and only if the fan $\Sigma$ is complete.
@@
If $J\neq\varnothing$, then $({\textbf{{v}}}\cdot{\textbf{{x}}}^{(n)})_{j}=0$ for any $j\in J$, ${\textbf{{v}}}\in V$ and $n=1,2,\ldots$ Consider the link subcomplex $\mathop{\mathrm{missing}}{link}_{\mathcal{K}}J=\{I\in\mathcal{K}\colon I\cup J\in\mathcal{K},\,I\cap J=\varnothing\}$. Then the data $\{\mathop{\mathrm{missing}}{link}_{\mathcal{K}}J,{\textbf{{a}}}_{j}\colon j\notin J\}$ define a complete fan in the quotient space $W^{*}/\langle{\textbf{{a}}}_{j}\colon j\in J\rangle$, which has dimension less than $W^{*}$. Therefore, we can apply the induction assumption in this case.
@@
The condition of completeness (or compactness of the quotient) can be formulated in terms of the triangulated configuration $(\mathcal{K},\mathrm{A})$ as follows.
@@
Assume that $\{\mathcal{K},\mathrm{A}\}$ is a triangulated configuration defining a simplicial fan $\Sigma$ in an $n$-dimensional space $W^{*}$. Then, the $\Sigma$ is complete if and only if one of the following two equivalent conditions is satisfied:
@@ 7. Real moment-angle manifolds
2. Let $(X,A)=(D^{1},S^{0})$, where $D^{1}$ is a closed interval $[-1,1]$ and $S^{0}$ is its boundary, consisting of two points. The polyhedral product $(D^{1},S^{0})^{\mathcal{K}}$ is known as the real moment-angle complex [13, §3.5][14]:
@@
We have an inclusion of pairs $(D^{1},S^{0})\hookrightarrow(\mathbb{R},\mathbb{R}^{\times})$, which induces an inclusion $\mathcal{R}_{\mathcal{K}}=(D^{1},S^{0})^{\mathcal{K}}\hookrightarrow(\mathbb{R},\mathbb{R}^{\times})^{\mathcal{K}}=U(\mathcal{K})$. So the required homeomorphism will follow from the fact that the $V$-orbit of any point ${\textbf{{x}}}=(x_{1},\ldots,x_{m})\in U(\mathcal{K})$ intersects $\mathcal{R}_{\mathcal{K}}$ at a single point.
@@
In [14, Theorem 4.7.5], it is proved that there is a homotopy equivalence $U(\mathcal{K})\simeq\mathcal{R}_{\mathcal{K}}$ for arbitrary $\mathcal{K}$. By virtue of Theorem 7.3, this homotopy equivalence can be realised as the quotient by a free action of $V\cong\mathbb{R}^{k}$ when $\mathcal{K}$ is the underlying complex of a complete simplicial fan.
@@ 8.1. Normal fans
is a fan in $W^{*}$, referred to as the normal fan of the polyhedron $P$. If $P$ is not full-dimensional, then the cones $\sigma_{Q}$ are not strictly convex, and $\Sigma_{P}$ is a generalised fan [16, Definition 6.2.2]
The support $|\Sigma_{P}|$ of a normal fan is convex (it coincides with the dual of the recession cone of $P$, see [16, Theorem 7.1.6]). In particular, $\Sigma_{P}$ is complete if and only if $P$ is a polytope.
@@
The normal fan $\Sigma_{P}$ of a polyhedron (8.1) consists of the cones $\mathop{\mathrm{cone}}\mathrm{A}_{I}$ such that $\sum_{i=1}^{m}b_{i}\gamma_{i}\in\mathop{\mathrm{relint}}\mathop{\mathrm{cone}}\Gamma_{\widehat{I}}$.
@@
where the second identity holds because $\Sigma$ has convex support. Since both $\Sigma_{P^{\prime}}$ and $\Sigma$ are subfans of $\Sigma_{P}$, we conclude that $\Sigma_{P^{\prime}}=\Sigma$. ∎
@@ 8.2. From polyhedra to quadrics
Now we turn to the quotient (leaf space) $U(\mathcal{K})/V$. As we have seen in Theorem 7.3, the quotient $U(\mathcal{K})/V$ is identified with the real moment-angle manifold $\mathcal{R}_{\mathcal{K}}$ when the data $\{\mathcal{K},{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ define a complete simplicial fan $\Sigma$. When $\Sigma$ is the normal fan of a polytope, there is another realisation of $U(\mathcal{K})/V$, which has an intrinsic smooth structure as a nondegenerate intersection of real quadratic surfaces.
@@
This result has well-known interpretations in the theory of Hamiltonian torus actions in symplectic geometry and in the theory of toric varieties (see [2, Proposition 3.1.1] and [23, Appendix 1, Theorems 1.2 and 1.4], and also §14.3 below). However, it is purely topological in nature.
@@
It remains to prove the key fact about the image of $L$. The argument below is based on the properties of the Legendre transform and is taken from [23, Appendix 1].
@@
This follows by combining Theorem 7.3 with Theorem 8.4 (4), as both $\mathcal{R}_{\mathcal{K}}$ and the intersection of quadrics are homeomorphic to the quotient $U(\mathcal{K})/V$ (see also [14, Theorem 6.2.4] for a direct argument). Note that the fan is assumed to be complete in Theorem 7.3, so we require that (8.1) is bounded, i. e. a polytope. ∎